# Magnetic Hysteresis Modeling with Neural Operators

### v1

@@ IV-B First-order reversal curves
The first neural operator utilized for modeling FORCs is the DeepONet. The operator uses 1000 random functions as input to the branch net and the $t_{\text{sample}}$ array of size 198 as input to the trunk net. The branch and trunk nets are fully connected feedforward neural networks, each comprising eight hidden layers with 200 neurons in each layer and a tanh activation function. The output dimension of both the branch and trunk nets is $p=100$. The model parameters are initialized using Xavier initialization [29], and the network is trained using the ADAM optimizer with a learning rate of $5\times 10^{-5}$ over 10000 epochs.
TABLE I: Errors in predicting first-order reversal curves ($\mathcal{R}$: relative error in $\mathbb{L}_{2}$ norm; MAE: mean absolute error; RMSE: root mean squared error.)
@@
The neural operators are subsequently benchmarked against traditional recurrent architectures, specifically RNN, LSTM, and GRU. Following the data processing similar to [13, 11], the input and output dimensions are set to 1000 for each cell. The sequence length is fixed at 198, corresponding to the size of the sampling array, $t_{\text{sample}}$. Each recurrent cell consists of 128 neurons with a single hidden layer. Training spans 10000 epochs using the ADAM optimizer with a learning rate $1\times 10^{-4}$.
Fig. 3 illustrates the input and output fields for training neural networks and corresponding results for the neural operators and sequential model in predicting magnetic hysteresis. The first row shows the input fields (blue curves) and output fields (black curves), along with predictions from four models: DeepONet, FNO, RIFNO, and LSTM, with FNO demonstrating the best approximation. The second row compares the FORC predictions for each model, where red curves indicate predictions and black curves show the ground truth. The third row presents absolute error plots, showcasing that LSTM performs poorly with widespread errors. The error for DeepONet propogates sinusoidally, FNO has a maximum error of 0.004, and the error in RIFNO predictions are prominent near the ends. Overall, neural operators perform better in approximating magnetic hysteresis than the recurrent neural methods.
Additionally, Table I compares the prediction errors for FORCs across six models: DeepONet, FNO, RIFNO, RNN, LSTM, and GRU, using three metrics: relative error ($\mathcal{R}$), MAE, and RMSE. FNO exhibits the lowest errors ($\mathcal{R}$ = 1.34e-3, MAE = 7.48e-4, RMSE = 9.74e-4), demonstrating superior accuracy. RIFNO also performs well with slightly higher errors ($\mathcal{R}$ = 1.63e-3, MAE = 8.79e-4, RMSE = 1.17e-3). DeepONet also captures the hysteresis well but with an order of magnitude less accuracy than FNO. In contrast, RNN, LSTM, and GRU show higher errors and cannot capture the hysteresis and predict for novel input $H$ fields. Table I underscores the effectiveness of FNO in accurately predicting first-order reversal curves, followed by RIFNO and DeepONet, while traditional recurrent models underperform.
IV-C Rate-independency
@@
TABLE II: Errors for testing on different rates
Rate Metric DeepONet FNO RIFNO
@@
MAE 6.37e-1 4.71e-2 8.79e-4
RMSE 8.14e-1 6.08e-2 1.17e-3
×$\times 10$ $\mathcal{R}$ 6.05e+0 9.05e-1 1.63e-3
MAE 3.71e+0 4.89e-1 8.79e-4
RMSE 4.37e+0 6.54e-1 1.17e-3
×$\times 100$ $\mathcal{R}$ 6.51e+0 2.40e+0 1.63e-3
MAE 4.20e+0 1.35e+0 8.79e-4
RMSE 4.70e+0 1.73e+0 1.17e-3
Figure 5: Performance of neural networks on predicting the minor loops. Top row: (From Left) the input $H$ fields (in blue) for the neural network, the output $B$ fields (in black), followed by the predictions (preds.) of the $B$ fields by DeepONet, FNO, RIFNO, and LSTM, respectively. Middle row: (From Left) the reference hysteresis loop (in black) followed by the predicted hysteresis loops (in red) compared with the reference (ref.) loops (in black) for the four methods. Bottom row: (From Left) absolute errors in predicting the minor loops for the four methods (in green).
@@
Fig. 4 shows the error metric $\mathcal{R}$ for DeepONet, FNO, and RIFNO across different testing rates compared to the training rate. The line graph demonstrates the variation of error for each model concerning the testing rate from $\frac{1}{100}$ to $100$ times the training rate. RIFNO exhibits the lowest and most consistent errors across all testing rates, indicating its superior generalization capability. FNO has higher errors compared to RIFNO. DeepONet shows the highest errors, especially at extreme testing rates, highlighting its comparatively lower robustness in varying conditions. Fig. 4 underscores the efficacy of RIFNO in maintaining low prediction errors across various testing scenarios, making it the most accurate operator among the three where different testing rates are considered.
Table II presents the errors in predicting the first-order reversal curve for different testing rates compared to the training rate using the DeepONet, FNO, and RIFNO models. The testing rates vary as $\frac{1}{100}$, $\frac{1}{10}$, $10$, and $100$ times the training rate, and the table includes three error metrics. For the $\frac{1}{100}$ rate, RIFNO demonstrates the lowest errors with $\mathcal{R}$ = 1.63e-3, MAE = 8.79e-4, and RMSE = 1.17e-3, outperforming DeepONet and FNO. This trend continues across all testing rates, with RIFNO consistently showing minimal errors, indicating its robustness and superior accuracy. In contrast, DeepONet exhibits the highest errors, especially at extreme testing rates. FNO performs better than DeepONet but still shows increasing errors as the testing rate deviates from the training rate.
IV-D Minor loops
TABLE III: Errors in predicting minor loops
@@ IV-D Minor loops
TABLE IV: Comparison of recurrent and operator architectures based on various attributes
@@
RNN LSTM GRU DON FNO RIFNO
@@
Maps functions to functions $\times$ $\times$ $\times$
Predictions on novel excitation fields $\times$ $\times$ $\times$
Rate-independent $\times$ $\times$
@@
Fig. 5 illustrates the performance of various neural operator methods in predicting the minor loops. The first row shows input fields (blue curves) and output fields (black curves), along with predictions from DeepONet, FNO, RIFNO, and LSTM, with RIFNO providing the most accurate approximations. The second row presents the minor loop curves, comparing predicted (red curves) and actual (black curves) output fields, demonstrating that FNO closely matches the ground truth. The third row displays absolute error plots, revealing that LSTM performs poorly with higher errors. The errors for DeepONet follow a specific pattern, while the errors for FNO predictions are minimal, with a maximum of approximately 0.02, and RIFNO errors are prominent near the ends.
Table III compares prediction errors for minor loops across six models: DeepONet, FNO, RIFNO, RNN, LSTM, and GRU, using relative error ($\mathcal{R}$), MAE, and RMSE. RIFNO achieves the lowest errors ($\mathcal{R}$ = 1.26e-2, MAE = 5.07e-3, RMSE = 7.03e-3), indicating superior accuracy. FNO and DeepONet follow with slightly higher errors than RIFNO. Conversely, traditional recurrent models like RNN, LSTM, and GRU show significantly higher errors, with LSTM and GRU performing the worst. Table III underscores the leading performance of RIFNO, followed by FNO and DeepONet, while traditional recurrent models lag in modeling novel minor loops not used for training the neural network model.
@@ V Discussion
A comparative analysis between recurrent and operator-based methods is presented in Table IV, focusing on essential attributes for modeling magnetic hysteresis. Table IV illustrates that operator learning methods excel in mapping arbitrary $H$ fields to $B$ fields while preserving the history of magnetization and demagnetization. The recurrent networks are rate-independent owing to their architecture, which does not explicitly include the rate at which the data is collected. RIFNO provides an advantage over recurrent and other operator networks by offering function-to-function mapping and rate independence properties.
@@
Future research directions include employing and developing different operator training methods. Techniques like transfer learning [34], leveraging pre-trained network parameters [35], and meta-learning approaches [36] can be utilized to accelerate model convergence. Additionally, the results presented herein for scalar hysteresis could be expanded upon to model vector hysteresis, leveraging the property that deep learning often circumvents the curse of dimensionality and generalizes well to higher dimensions.
@@ VI Conclusions
This paper introduced a novel approach to model magnetic hysteresis through deep learning. The research highlights the benefits of employing neural operators for hysteresis modeling, yielding generalizable models essential for scenarios where prior training on varying magnetic fields is impractical. By employing prominent operator networks, namely deep operator network and Fourier neural operator, the research demonstrated the prediction of novel first-order reversal curves and minor loops outperforming the traditional recurrent architectures and leading to the development of generalizable neural hysteresis models. Furthermore, this contribution introduced a rate-independent Fourier neural operator, which enhanced the neural operator framework by incorporating the rate-independent hysteresis characteristics prevalent in magnetic materials. Concretely, this paper advances neural hysteresis modeling by developing accurate and generalizable models capable of real-time inference of material responses. The codes and dataset related to the work will be made available upon publication.

### v2

@@ IV-C First-order reversal curves
The first neural operator utilized for modeling FORCs is the DeepONet. The operator uses 1000 random functions as input to the branch net and the $t_{\text{sample}}$ array of size 198 as input to the trunk net. The branch and trunk nets are fully connected feedforward neural networks, each comprising eight hidden layers with 200 neurons in each layer and a tanh activation function. The output dimension of both the branch and trunk nets is $p=25$, following the hyperparameter selection experiment. The model parameters are initialized using Xavier initialization [48], and the network is trained using the ADAM optimizer with a learning rate of $5\times 10^{-5}$ over 10000 epochs.
TABLE IV: Errors in predicting first-order reversal curves ($\mathcal{R}$: relative error in $\mathbb{L}_{2}$ norm; MAE: mean absolute error; RMSE: root mean squared error.)
@@
EDLSTM 1.64e+0 9.81e-1 1.18e+0
Furthermore, WNO is employed to predict FORCs. WNO inputs are identical to FNO, and are concatenated data consisting of 1000 instances of $H$ fields and a sampling array repeated across 1000 samples. An initial fully connected layer is utilized to project inputs from 2 channels to 64 channels, i.e., $N_{\text{f}}=64$ , followed by eight wavelet integral layers ($L=8$), each with 64 input and 64 output channels. Within each block, four levels of wavelet decomposition ($n_{\text{l}}=4$) are performed employing db6 wavelet. The networks conclude with two fully connected layers transforming dimensions from 64 channels to 128 and finally to a single output channel. GeLU activation function is applied throughout the network. Training parameters encompass a batch size of 100 samples over 10000 epochs, utilizing the ADAM optimizer with a learning rate of $1\times 10^{-3}$.
The neural operators are subsequently benchmarked against traditional recurrent architectures, specifically RNN, LSTM, GRU, and recently proposed encoder-decoder LSTM (EDLSTM) [14]. Following the data processing similar to [14, 12], the input and output dimensions are set to 1000 for each cell. The sequence length is fixed at 198, corresponding to the size of the sampling array, $t_{\text{sample}}$. Each recurrent cell consists of 128 neurons with a single hidden layer. Training spans 10000 epochs using the ADAM optimizer with a learning rate $1\times 10^{-4}$.
@@
Furthermore, neural operators are compared with EDLSTM, a recurrent neural architecture to model magnetic hysteresis, as presented in [14]. The architecture comprises an encoder and a decoder. The encoder consists of a 1-layer 128-state LSTM network, processing the input sequence by leveraging the LSTM gate mechanisms, allowing the encoder to capture and store the characteristics of the input sequence as hidden and cell states. The decoder mirrors the architecture of the encoder. It generates the output sequence through an autoregressive inference method, iteratively updating by feeding the previous output back into the network to predict the next time step until the entire output sequence is generated. Training is carried out for 10000 epochs using the ADAM optimizer with a learning rate $1\times 10^{-4}$ as in the case of LSTM.
Fig. 4 illustrates the results for the neural operators and sequential model in predicting magnetic hysteresis. The first row shows the predictions from six models: DeepONet, FNO, WNO, RIFNO, LSTM, and EDLSTM, with FNO demonstrating the best approximation. The second row compares the FORC predictions for each model, where red curves indicate predictions and black curves show the ground truth. The third row presents absolute error plots, showcasing that LSTM, and EDLSTM perform poorly with widespread errors. The error for DeepONet propogates sinusoidally, FNO and WNO have a maximum error around 0.004 and 0.01, respectively, and the error in RIFNO predictions are prominent near the ends. Overall, neural operators perform better in approximating magnetic hysteresis than the recurrent neural methods.
Additionally, Table IV compares the prediction errors for FORCs across eight models: DeepONet, FNO, WNO, RIFNO, RNN, LSTM, GRU, and EDLSTM using three metrics: relative error ($\mathcal{R}$), MAE, and RMSE. FNO exhibits the lowest errors ($\mathcal{R}$ = 1.34e-3, MAE = 7.48e-4, RMSE = 9.74e-4), demonstrating superior accuracy. RIFNO also performs well with slightly higher errors ($\mathcal{R}$ = 1.63e-3, MAE = 8.79e-4, RMSE = 1.17e-3). DeepONet and WNO also capture the hysteresis well but with an order of magnitude less accuracy than FNO. In contrast, RNN, LSTM, GRU, and EDLSTM show higher errors and cannot capture the hysteresis and predict for novel input $H$ fields. Table IV underscores the effectiveness of FNO in accurately predicting first-order reversal curves, followed by RIFNO, DeepONet, and WNO while traditional recurrent models underperform.
IV-D Rate-independency
TABLE V: Errors for testing on different rates
Rate Metric DeepONet FNO WNO RIFNO
@@ IV-D Rate-independency
MAE 6.37e-1 4.71e-2 5.74e-1 8.79e-4
RMSE 8.14e-1 6.08e-2 9.08e-1 1.17e-3
×$\times 10$ $\mathcal{R}$ 6.05e+0 9.05e-1 8.72e+0 1.63e-3
MAE 3.71e+0 4.89e-1 3.27e+0 8.79e-4
RMSE 4.37e+0 6.54e-1 9.49e+0 1.17e-3
×$\times 100$ $\mathcal{R}$ 6.51e+0 2.40e+0 9.51e+0 1.63e-3
MAE 4.20e+0 1.35e+0 4.12e+0 8.79e-4
RMSE 4.70e+0 1.73e+0 5.33e+0 1.17e-3
Figure 6: Performance of neural networks on predicting the minor loops. Top row: (From Left) the predictions (preds.) of the $B$ fields by DeepONet, FNO, WNO, RIFNO, LSTM, and Encoder-Decoder LSTM (EDLSTM) respectively. Middle row: (From Left) the predicted hysteresis loops (in red) compared with the reference (ref.) loops (in black) for the six methods. Bottom row: (From Left) absolute errors in predicting the minor loops for the six methods (in green).
@@
Fig. 5 shows the error metric $\mathcal{R}$ for DeepONet, FNO, RIFNO, and WNO across different testing rates compared to the training rate. The line graph demonstrates the variation of error for each model concerning the testing rate from $\frac{1}{100}$ to $100$ times the training rate. RIFNO exhibits the lowest and most consistent errors across all testing rates, indicating its superior generalization capability. FNO has higher errors compared to RIFNO. DeepONet, and WNO show the highest errors, especially at extreme testing rates, highlighting their comparatively lower robustness in varying conditions. Fig. 5 underscores the efficacy of RIFNO in maintaining low prediction errors across various testing scenarios, making it the most accurate operator among the three where different testing rates are considered.
Table V presents the errors in predicting the first-order reversal curve for different testing rates compared to the training rate using the DeepONet, FNO, RIFNO, and WNO models. The testing rates vary as $\frac{1}{100}$, $\frac{1}{10}$, $10$, and $100$ times the training rate, and the table includes three error metrics. For the $\frac{1}{100}$ rate, RIFNO demonstrates the lowest errors with $\mathcal{R}$ = 1.63e-3, MAE = 8.79e-4, and RMSE = 1.17e-3, outperforming DeepONet, FNO, and WNO. This trend continues across all testing rates, with RIFNO consistently showing minimal errors, indicating its robustness and superior accuracy. In contrast, DeepONet, and WNO exhibiting the highest errors, especially at extreme testing rates. FNO performs better than DeepONet, and WNO but still shows increasing errors as the testing rate deviates from the training rate.
IV-E Minor loops
TABLE VI: Errors in predicting minor loops
@@ IV-E Minor loops
EDLSTM 1.01e+0 4.68e-1 5.61e-1
TABLE VII: Comparison of recurrent and operator architectures based on various attributes
@@
RNN LSTM GRU EDLSTM DeepONet FNO WNO RIFNO
@@
Maps functions to functions $\times$ $\times$ $\times$ $\times$
Predictions on novel excitation fields $\times$ $\times$ $\times$ $\times$
Rate-independent $\times$ $\times$ $\times$
@@
Fig. 6 illustrates the performance of various neural operator methods in predicting the minor loops. The first row shows predictions from DeepONet, FNO, WNO, RIFNO, LSTM, and EDLSTM with RIFNO providing the most accurate approximations. The second row presents the minor loop curves, comparing predicted (red curves) and actual (black curves) output fields, demonstrating that FNO closely matches the ground truth. The third row displays absolute error plots, showing that LSTM, and EDLSTM perform poorly with higher errors. The errors for DeepONet follow a specific pattern, while the errors for FNO predictions are minimal, with a maximum of approximately 0.02, whereas WNO predictions have a maximum absolute error around 0.01. The errors for RIFNO are prominent near the ends.
Table VI compares prediction errors for minor loops across eight models: DeepONet, FNO, WNO RIFNO, RNN, LSTM, GRU, and EDLSTM using relative error ($\mathcal{R}$), MAE, and RMSE. RIFNO achieves the lowest errors ($\mathcal{R}$ = 1.26e-2, MAE = 5.07e-3, RMSE = 7.03e-3), indicating superior accuracy. FNO and DeepONet follow with slightly higher errors than RIFNO. WNO exhibits the maximum error amongst the neural operator methods. Conversely, traditional recurrent models like RNN, LSTM, GRU and recent methods like EDLSTM show significantly higher errors, with RNN performing the worst. Table VI underscores the leading performance of RIFNO, followed by FNO, DeepONet, and WNO while traditional recurrent models lag in modeling novel minor loops not used for training the neural network model.
@@ V Discussion
A comparative analysis between recurrent and operator-based methods is presented in Table VII, focusing on essential attributes for modeling magnetic hysteresis. Table VII illustrates that operator learning methods excel in mapping arbitrary $H$ fields to $B$ fields while preserving the history of magnetization and demagnetization. The recurrent networks are rate-independent owing to their architecture, which does not explicitly include the rate at which the data is collected. RIFNO provides an advantage over recurrent and other operator networks by offering function-to-function mapping and rate independence properties.
@@
Future research directions include employing and developing different operator training methods. Techniques like transfer learning [53], leveraging pre-trained network parameters [54], and meta-learning approaches [55] can be utilized to accelerate model convergence. Additionally, the results presented herein for scalar hysteresis could be expanded upon to model vector hysteresis, leveraging the property that deep learning often circumvents the curse of dimensionality and generalizes well to higher dimensions. Furthermore, validating neural operators for experimental data is part of our future work, where the study will be extended to include experimental data for different magnetic materials.
@@ VI Conclusions
This paper introduced a novel approach to model magnetic hysteresis through neural operators. The research highlights the benefits of employing neural operators for hysteresis modeling, yielding generalizable models essential for scenarios where prior training on varying magnetic fields is impractical. By employing prominent operator networks, namely deep operator network, Fourier neural operator, and wavelet neural operator, the research demonstrated the prediction of novel first-order reversal curves and minor loops outperforming the traditional recurrent architectures and leading to the development of generalizable neural hysteresis models. Furthermore, this contribution introduced a rate-independent Fourier neural operator, which enhanced the neural operator framework by incorporating the rate-independent hysteresis characteristics prevalent in magnetic materials. Concretely, this paper advances neural hysteresis modeling by developing accurate and generalizable models capable of real-time inference of material responses.